### v3

# Evaluating LLMs on Entity Disambiguation in Tables

Abstract.
Tables are crucial containers of information, but understanding their meaning may be challenging. Over the years, there has been a surge in interest in data-driven approaches based on deep learning that have increasingly been combined with heuristic-based ones. In the last period, the advent of Large Language Models (LLMs) has led to a new category of approaches for table annotation. However, these approaches have not been consistently evaluated on a common ground, making evaluation and comparison difficult. This work proposes an extensive evaluation of four STI state-of-the-art (SOTA) approaches — Alligator (formerly s-elBat), Dagobah, TURL, and TableLlama; the first two belong to the family of heuristic-based algorithms, while the others are respectively encoder-only and decoder-only LLMs. We also include in the evaluation both GPT-4o and GPT-4o-mini, since they excel in various public benchmarks. The primary objective is to measure the ability of these approaches to solve the entity disambiguation task with respect to both the performance achieved on a common-ground evaluation setting and the computational and cost requirements involved, with the ultimate aim of charting new research paths in the field.
Semantic Web, Knowledge Base, Knowledge Base Construction, Knowledge Base Extension, Knowledge Graph, Semantic Table Interpretation, Table Annotation, Data Enrichment, Tabular Data
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2018; Boise, Idaho, USAccs: Computing methodologies Machine learningccs: Information systems Ontologies
1. Introduction
Tables are commonly used to create, organise, and share information in various knowledge-intensive processes and applications in business and science. Disambiguating values occurring in the table cells using a background Knowledge Graph (KG) is useful in different applications. First, it is part of the broader objective of letting machines understand the table content, which has been addressed by different research communities with slightly different formulations like Semantic Table Interpretation (STI(Liu2023, ) - the one considered in this paper, semantic labelling (pham2016semantic, ), and table annotation (Suhara2022, ). The main idea behind these efforts is to match the table against a background knowledge graph by annotating cells (mentions) with entities (Cell-Entity Annotation - CEA), columns with class labels (Column-Type Annotation - CTA), and pairs of columns with properties (Column-Property Annotation - CPA) (Jimenez2020, ). Second, annotations produced by table-to-graph matching algorithms can be used to transform the tables into Knowledge Graphs (KGs) or populate existing ones. Third, links from cells to entities of KGs support data enrichment processes by serving as bridges to augment the table content with additional information (Cremaschi2020-1, ). This conceptualisation covers most of the proposed definitions by generalizing some aspects (e.g., consideration of NILs and selection of column pairs to annotate).
In particular, we focus on Entity Linking (EL) in tables (CEA, in the STI terminology), which is relevant not only to support table understanding but also to support data transformation, integration and enrichment processes, which are particularly interesting from a data management point of view. The Cell-Entity Annotation (CEA) task can be broken down into two sub-tasks: candidate Entity Retrieval (ER), where a set of candidates for each mention is collected and, often, associated with an initial score and rank, and Entity Disambiguation (ED), where the best candidate is selected (and, in some case, a decision whether to link or not is also considered (Avogadro2023, )). When considering approaches to STI and, especially, CEA, it should be considered that the content and the structure of tables may differ significantly, also depending on application-specific features: column headers may have interpretable labels or be omitted; the number of rows can vary from a dozen (e.g., as typical in tables published on the web or in scientific papers) to hundreds thousand or even millions (e.g., in business data); the cells may include reference to well-known entities (e.g., geographical entities) as well as to specific ones (e.g., biological taxa); tables may come with a rich textual context (e.g., caption or other descriptions in web or scientific documents), or no context at all (e.g., in business data) (Liu2023, ).
A first generation of approaches for CEA has exploited different matching heuristics, traditional machine learning approaches based on engineered features (in the following “feature-based ML”), or a combination of both (Liu2023, ).
Figure 1. Architectures of TableLlama, TURL and Alligator.
[…]
Architectures of TableLlama, TURL and Alligator. Refer to caption
The increased recognition of the power of LLMs has led to a new generation of Generalistic Table Understanding and Manipulation (GTUM) approaches that support STI (CEA, CTA, and CPA) and other tasks (e.g., question answering, schema augmentation, row population, and fact verification among the others). The first remarkable example is TURL (Deng2022, ), which is based on an adaptation of BERT (Devlin2019, ) to consider the tabular structure, additionally fine-tuned to execute specific tasks (including ED). The latest of these models, TableLlama (Zhang2023, ), is based on the autoregressive Large Language Model (LLM) Llama 2-7B (Touvron2023-2, ), which is fine-tuned with instruction tuning(wei2022instruction, ) to perform specific tasks. These approaches have been trained and tested on datasets that fit their generalistic ambition and reported as SOTA approaches for the considered tasks.
While both STI-related and GTUM approaches facilitate entity disambiguation (ED), a comparative evaluation of these two method families on a common experimental foundation has yet to be conducted. STI-related methods have primarily been assessed within the International Semantic Web Challenge on Tabular Data to Knowledge Graph Matching (SemTab), where most tables are synthetically generated by querying knowledge graphs (KGs), closely resembling tables from industrial or corporate environments. These tables, however, lack the rich semantic context of web tables found on platforms like Wikipedia. In contrast, GTUM approaches have been evaluated mainly on web-based tables, which are semantically rich and crafted by humans for human use. Consequently, it remains unclear how these approaches perform relative to the state of the art (SOTA) under different conditions. Finally, several aspects of CEA that are important from a data management point of view, e.g., scalability, have not been considered. We posit that several of these questions are important to understand the current status of ED algorithms for tables, as well as the challenges to be addressed in the future.
With our study, we want to provide a more detailed analysis of the advantages and disadvantages of GTUM ED approaches based on LLMs, especially when compared to those specifically focusing on CEA developed in the context of STI. Our analysis mainly considers the SOTA generalistic generative model TableLlama (Zhang2023, ), its predecessor TURL (Deng2022, ), a BERT-based encoder-only transformer, and Alligator (Avogadro2023, ), a recent feature-based Machine Learning (ML) approach. These models represent three different inference mechanisms for ED, sketched in Figure 1, each one associated with a set of expected advantages: TableLlama is expected to exploit implicit knowledge of a large LLM; TURL is expected to be a more efficient generalistic model based on a small LLM fine-tuned specifically for the ED task; Alligator exploits a set of features engineered based on the experience with the SemTab challenge, which is further processed by two small neural networks to return confidence scores. Each approach is tested in in-domain, out-of-domain, and moderately out-of-domain settings, as more precisely defined in Section 4.2. In addition, we test the moderately out-of-domain fine-tuning to test generalisation and adaptation capabilities of TURL, TableLlama and Alligator. Finally, we provide the results obtained by Dagobah, an approach that achieved top performance in previous SemTab challenges [^1] With some limitations due to its excessive execution time and the difficulty of properly replicating the published results despite our best effort..
We remark that this paper aims to provide insights on the behavior and impact of GTUM models on ED rather than introducing a new approach. This is inspired by the large body of similar analyses addressing specific problems in NLP (Wang2023, ; kirk2024understanding, ; zhong2024can, ; zhang2024llmeval, ; kandpal2023llmstruggle, ). In our work, we also develop an evaluation protocol that can be used in future research.
To summarise, our main contributions are:
  1. (1)
    Test the performance of different genres of STI models when used in combination with a realistic candidate retrieval method;
  2. (2)
    Test the performance of these models on several datasets used to evaluate SOTA STI approaches, through a common-ground comparison with approaches trained on these data;
  3. (3)
    Assess the performance improvement by adaptation with additional moderately-out-of-domain fine-tuning (Sec. 4.2);
  4. (4)
    Evaluate the computational efficiency and provide implications for actual usage in different application settings.
This paper is organised as follows: Section 2 proposes a detailed examination of the techniques used by STI approaches in the SOTA. Section 3 introduces and details the approaches tested in this work, relating them to the ED challenges they are intended to solve. Section 4 describes the objectives of this study, the datasets used to evaluate the selected approaches and defines the experimental settings followed in Section 5, which introduces the configuration parameters and the evaluation results, discussing the main results and the ablation studies. Finally, we conclude this paper and discuss the future direction in Section 6.
2. Related Work
CEA is usually divided into two sub-tasks: retrieval of candidate entities, i.e., ER (Ratinov2009, ), and entity disambiguation, i.e., ED (Rao2013, ), where one or no candidate is selected as a link, usually after scoring and ranking the candidates. Approaches to ED for tables have been mostly proposed as part of broader STI systems (Liu2023, ). For this paper, we group these approaches into two main categories: i) those based on heuristics, (feature-based) ML and probabilistic approaches, and ii) those based on LLM.
Heuristic, ML and Probabilistic-based Approaches. We refer to (Liu2023, ) for a thorough review of STI approaches and ED approaches proposed therein up to 2022. The ED task in STI can be performed by applying multiple techniques while focusing on different inherent information: i) similarity, ii) contextual information, iii) ML techniques, and iv) probabilistic models. Often, the disambiguation step involves selecting the winning candidate based on heuristics, like the string similarity between the entity label and mention (Limaye2010, ; Zhang2017, ; Oliveira2019, ; Kruit2019, ; Steenwinckel2019, ; Thawani2019, ; Cremaschi2019, ; Cremaschi2020-1, ; Cremaschi2020-2, ; Huynh2020, ; Tyagi2020, ; Shigapov2020, ; Avogadro2021, ; Abdelmageed2021, ; Huynh2021, ; Chen2022, ; Cremaschi2022, ).
Contextual information during the CEA task considers the surrounding context of a table cell, such as neighbouring cells, column headers, or header row. Contextual information provides additional clues or hints about the meaning and intent of the mention. By analysing the context, a system can better understand the semantics of the cell and make more accurate annotations (Syed2010, ; Ritze2015, ; Efthymiou2017, ; Chen2019, ; Nguyen2019, ; Thawani2019, ; Nguyen2020, ; Nguyen2021, ; Morikawa2019, ; Azzi2020, ; Chen2020, ; Cremaschi2020-2, ; Abdelmageed2020, ; Huynh2020, ; Avogadro2021, ; Baazouzi2021, ; Huynh2021, ; Huynh2022, ; Chen2022, ; Cremaschi2022, ; Liu2022, ).
Other methods that can be employed are ML techniques. These techniques typically involve training a ML model on a labelled dataset where cells are annotated with their corresponding entities. The models, like Support Vector Machine (SVM) (Mulwad2010, ), Neural Network (NN) (Thawani2019, ) and Random Forest (Zhang2020, ), learn patterns and relationships between the cells content and their associated entities. To predict the most appropriate entity, ML techniques consider various cell features, such as the textual content, context, neighbouring cells, and other relevant information. Alligator is a recent approach belonging to this family (Avogadro2023, ); it is used in this study and further described in Section 3.
Probabilistic models are frameworks for representing and reasoning under uncertainty using probability theory. These models vary in their representation of dependencies and use diverse graphical structures. Several Probabilistic Graphical Model (PGM) can also be used to resolve the disambiguation task, such as Markov models or Loopy Belief Propagation (LBP) (Mulwad2011, ; Mulwad2013, ; Bhagavatula2015, ; Kruit2019, ; Yang2021, ).
LLMs-based approaches. In the current SOTA, several attempts to apply LLM in the STI process can be identified. Based on the architecture structure of LLMs, approaches can be categorised into two groups: i) encoder-based, and ii) decoder-based (Pan2024, ).
Starting from encoder-based approaches, Ditto (Li2020, ) utilizes Transformer-based language models to perform a slightly different task, i.e., entity-matching between different tables. Doduo (Suhara2022, ) performs Column-Type Annotation (CTA) and CPA using a pre-trained language model, precisely fine-tuning a BERT model on serialised tabular data. Column (columns-properties) types are predicted using a dense layer followed by an output layer with a size corresponding to the number of column (columns-properties) types. Dagobah SL 2022 (Huynh2022, ) employs an ELECTRA-based (electra2020, ) cross-encoder, a variant of the BERT model, which takes a concatenated input, including table headers and the entity description, and outputs a probability value representing the entity’s likelihood concerning the headers. TorchicTab (Dasoulas2023, ) is composed of two sub-systems: TorchicTab-Heuristic and TorchicTab-Classification, with the latter that utilises Doduo (Suhara2022, ) internally.
Regarding the decoder-only approaches, (Korini2023, ) explored the CTA task by employing ChatGPT, performing experiments with diverse prompts tailored for the task using a subset of the SOTAB benchmark (Korini2022, ). Another study evaluates GPT3.5-turbo-0301 in zero-shot settings on a task that was somehow related to CEA; the task consisted of classifying descriptions of products based on attribute-value pairs (Peeters2023, ). Some works included structured tabular data into the training process of general purpose decoder-based LLMs to address the peculiarities of specific domains, e.g., BloombergGPT (Wu2023, ) for the financial domain or CancerGPT for the medical one (li2024cancergpt, ). Other works, instead, focused on table-specific tasks, especially reasoning (suiWSDM2024, ; kong2024opentab, ; nahid2024tabsqlifyenhancingreasoningcapabilities, ), enhanced by Chain-of-Thought (CoT) (wang2024chainoftable, ; zheng-etal-2023-chain, ). Another decoder-based approach, TableGPT (Li2023, ), performs several tasks, including CTA using a fine-tuned version of GPT-3.5.
Of particular interest for our work, LLM-based models capable of addressing all the STI tasks and in particular the CEA one are: i) TURL (Deng2022, ), which leverages a pre-trained TinyBERT (Jiao2020, ) model to initialise a structure-aware Transformer encoder, fine-tuned to obtain contextualised representations for each cell, with matching scores between the KG candidates’ representations and cell embeddings that are calculated using a linear function, and transformed into a probability distribution over the candidate entities; ii) TableLlama (Zhang2023, ), a fine-tuned Llama2 on a multi-task dataset for tabular data, performs CEA ( along with several other tasks), where the entity linking sub-dataset derives from the TURL (Deng2022, ) dataset, by prompting the LLM to retrieve the correct candidate given the serialized table, the table metadata, the cell to be linked and a set of proposed candidates.
A more comprehensive review of both encoder-only and decoder-only LLM-based approaches for Table Understanding can be found in (fang2024llm-table-survey, ).
Table 1. Statistics of the datasets used to pre-train the selected approaches
Approach Dataset Entities Source Entities Domain
Alligator (Avogadro2023, ) SemTab2021 - R2 47.4K Graph Queries Cross
SemTab2021 - R3 58.9K Graph Queries
SemTab2022 - R2 (2T) 994.9K Graph Queries / Web Tables
SemTab2020 - R4 667.2K Graph Queries
SemTab2019 - R3 390.4K Graph Queries
SemTab2019 - R1 (T2D) 8K Graph Queries / Web Tables
Total 2.16M
TURL (Deng2022, ) TURL WikiTable w/o WikiGS (efthymiou2017wikigs, ) 1.23M Web Tables Cross
TableLlama (Zhang2023, ) TableInstruct (Zhang2023, ) 2.6M Web Tables Cross
3. Selected Approaches
We select four STI approaches representative of different categories of algorithms that solve semantic tasks on tables, and, especially CEA: TableLlama is the first autoregressive LLM that is specifically instruction-tuned on tabular data and reports SOTA results (Zhang2023, ); TURL, cited as previous SOTA in (Zhang2023, ), is a BERT-based encoder-only model performing the three STI tasks, including CEA (Deng2022, ); Alligator is a recent feature-based ML algorithm focusing on CEA, is publicly available, and has been evaluated in settings similar to the moderately-out-of-domain settings discussed in this paper (Avogadro2023, ); Dagobah, a heuristic-based algorithm, has been the winner of various rounds of the SemTab challenge in 2020 (Huynh2020, ), 2021 (Huynh2021, ) and 2022 (Huynh2022, ) and is publicly available.
While we focus our study only on the CEA task, we remark that all the selected approaches perform all three STI tasks (which is rare and not obvious, especially for the LLM-based approaches), i.e., CTA, CPA and CEA.
TURL (Deng2022, ) introduces the standard BERT pre-training/finetuning paradigm for relational Web tables and is composed of three main modules: i) an embedding layer to convert different components of a table into embeddings, ii) a pre-trained TinyBERT Transformer (Jiao2020, ) with structure-awareness to capture both the textual information and relational knowledge, and iii) a final projection layer for pre-training/fine-tuning objectives. The embedding layer is responsible for embedding the entire table, distinguishing between word embeddings $\mathbf{x}_{w}$, derived from table metadata and cell text (mentions), and entity embeddings $\mathbf{x}_{e}$, representing the unique entities to be linked. The sequence of tokens in $\mathbf{x}_{w}$ and the entity representation $\mathbf{x}_{e}$ are sent to a structure-aware TinyBERT Transformer to obtain a contextualised representation. To inject the table structural information into the contextualised representations, a so-called visibility-matrix $M$ is created, such that entities and text content in the same row or column are visible to each other, except table metadata that are visible to all table components. During pre-training, both the standard Masked Language Model (MLM) and the newly introduced Masked Entity Retrieval (MER) objectives are employed, with the projection layer that is learned to retrieve both the masked tokens and entities [^2] The entities are retrieved from a small candidates set, considering that entity vocabulary could be quite large. During fine-tuning, the model is specialised to address the specific downstream task. Specifically, for CEA, when provided with the sub-table containing all mentions to be linked to an external KG, TURL is fine-tuned to produce a probability distribution over the combined set of candidates for each mention to be linked. This means that TURL lacks the context provided by all the not-to-be-linked cells, without the possibility of abstaining from answering or responding with a NIL. Also, in the original paper, we observe that the best candidate is chosen by a function that compares the best score computed by TURL and the score assigned by the model to the first candidate retrieved by the Wikidata-Lookup service by down-weighting the best model prediction. As in previous comparisons (Zhang2023, ), to evaluate TURL’s performance on the ED task, we consider the model predictions only. Finally, as TURL accepts the whole table as input, computational performance varies depending on the size of the input table.
TableLlama (Zhang2023, ) employs Llama2 (Touvron2023-2, ) and instruction tuning (wei2022instruction, ) to solve multiple tasks related to tables, e.g., CTA, CPA, CEA and Q&A to name a few. To this end a multi-task instruction tuning dataset named TableInstruct is made available, containing more than 1.24M training tables and 2.6M instances gathered from 14 different table datasets of 11 distinctive tasks. To account for longer tables, exceeding the maximum context length of Llama2 (which is fixed at 4096 tokens), LongLora (Chen2024, ) is used to enlarge the maximum context length to 8192 tokens. In particular, the LLM is prompted with i) an instruction that describes the high-level task to be solved, ii) an input prepended by the [TLE] special character followed by the table metadata, if available, and the serialised table and iii) a question based on the task to be solved by the LLM. In particular, the table starts with the [TAB] special character and is followed first by the table header and then by every row in the table, separated by the special character separator [SEP]. For the CEA task, the question asks the LLM to link a particular mention found in the table against a small set of candidates (maximum 50 candidates) by extracting the correct candidate from the proposed list without, as in TURL, the possibility of not answer or to answer with a NIL. Since the context is fixed to 8192 tokens, compromises must be made to reduce either the length of the set of candidates or the input table.
Table 2. Comparison of TURL, TableLlama, and Alligator for the CEA task.
TURL (Deng2022, ) TableLlama (Zhang2023, ) Alligator (Avogadro2023, )
Model TinyBERT (Jiao2020, ) equipped with table structure-awareness (visibility matrix) Llama2 with LongLora (Chen2024, ) for extended context Feature-based ML approach with two Neural Network (NN)s for scoring
Input Table comprised of only the mentions to be linked, optional metadata and candidates for every mentions Serialized table with instruction, optional metadata, mention to be linked and the list of candidates, limited to 8192 tokens Whole table (candidates are retrieved on-the-fly thanks to LamAPI)
Objective Pre-training with MLM and MER; task-specific fine-tuning on CEA Multi-task instruction tuning for CEA, CTA, CPA, and other table-related tasks Feature-based binary cross-entropy with negatives sampling: ${p(y_{m}=1|F_{m})\geq p(y_{m}=0|F_{w};w\in\mathcal{N}(m))}$ [^3] Given a mention $m$ and its correct associated entity in the knowledge graph, $p(y_{m}=1\mid F_{m})$ represents the probability that the features $F_{m}$ correctly identify the correct candidate. For each incorrect candidate $w$ in the set $\mathcal{N}(m)$, $p(y_{m}=0\mid F_{w})$ represents the probability that the features $F_{w}$ correctly indicate that $w$ is not the correct candidate.
Output Probability distribution over candidates (Auto-regressively) Extracts candidate directly from list Score in $[0,1]$ for every candidate
Limitations Input table limited to mentions to be linked, no NIL-awareness Limited to 8192 tokens, no score over candidates, no NIL-awareness No semantic-related features, biased towards KG matching
Alligator is a feature-based ML approach based on multiple steps: i) data analysis and pre-processing, ii) ER, iii) local feature extraction, iv) local scoring, v) feature enrichment, vi) context-based scoring. The data analysis and pre-processing step applies standard pre-processing transformations (e.g., lowercase, space removing), with columns classified as either L-column (containing literals) or NE-column (containing named-entity mentions). The ER step extracts relevant candidates from the KG using the LamAPI services. The local feature extraction step builds a vector of engineered features for each candidate entity; the vector represents information about the specific candidate (e.g., its KG popularity, number of tokens and the syntactical similarity between the candidate label and the mention) or its comparison with the row cells content (e.g., trying to match related literals with the row content). Local scoring computes a matching score for each candidate, using a simple deep NN (denoted PN in (Avogadro2023, )): the NN takes the local feature vectors as input and is trained to convert the features into a normalised matching score. In practice, this step re-ranks all the candidates, considering mainly text similarity and matches against values on the same row. The last steps compute updated scores for the candidate of a given mention by considering the best candidates for other cells on the same column. In practice, Columns-Property Annotation (CPA) and CTA tasks are executed on a best-effort basis to capture, for each candidate, the degree of agreement between its types and properties and the types of properties of other best candidates in the same column. After this feature enrichment step, a simple deep NN (denoted RN in (Avogadro2023, )) similar to the previous one predicts the updated normalised scores for all the candidates [^4] In principle, the approach also estimates a confidence score for each cell to make a decision whether to link the best candidate or not; however, in this paper, we focus on the ED task and consider the candidate with the best score as the output of Alligator.
Dagobah (Huynh2021, ) performs the CEA, CPA and CTA tasks by implementing a multi-step pipeline: i) table pre-processing, ii) ER, iii) candidate pre-scoring, iv) CPA, CTA, and v) CEA. The table pre-processing step involves estimating table metadata, e.g., header detection, and column type estimation. These tasks help identify the structure and annotation targets, facilitating subsequent annotation steps. ER retrieves relevant candidate entities from the KG for each mention in the table using Elasticsearch; it considers the primitive types identified during pre-processing and enriches entity information with aliases to increase coverage. Then, the candidate pre-scoring computes a relevance score for each candidate by combining two features: context and literal similarity. Given a mention to disambiguate and a candidate entity, the context feature compares the mention neighbors to the candidate neighbors with respect to the underline KG. Finally, the most relevant entity for each mention is selected by combining pre-scoring with information from CPA and CTA into a final score. The main challenges with Dagobah are its high time complexity and memory usage, making it impractical for processing large tables. This difficulty stems from the expensive task of matching table rows with relations in a KG. Consequently, we limited our evaluation of Dagobah to the datasets HT-R1, HT-R2, and WikidataTables-R1 because these datasets contain tables with limited size, as detailed in Section 4.1.
Table 1 lists the datasets used to pre-train the selected approaches; in particular, for TableLlama we use the published model on HugginFace, while we train TURL and Alligator on the same datasets used in the original paper. We use the term ”pre-train” to refer to this main training phase to distinguish it from subsequent fine-tuning (cfr. Section 4.4). Dagobah is not listed because is not based on training.
Table 3. Statistics of the datasets used to fine-tune and evaluate models. †indicates datasets that have been sub-sampled so that they can be entirely processed by TableLlama within its 8192 context. {Dataset}-red means that the specific dataset has been reduced as explained in Section 4.1.
Dataset Split Tables Cols (min |||| max |||| $\bar{x}$) Rows (min |||| max |||| $\bar{x}$) Entities
Train† 91 1 |||| 8 |||| 4,86 5 |||| 369 |||| 98,61 14 674
SemTab2022 2T-red Test† 26 1 |||| 8 |||| 4,65 7 |||| 264 |||| 74,58 4 691
Train 3 691 2 |||| 5 |||| 2,56 4 |||| 8 |||| 5,68 26 189
SemTab2022 HT-R1 Test 200 2 |||| 5 |||| 2,59 4 |||| 8 |||| 5,74 1 406
Train† 4 344 2 |||| 5 |||| 2,56 4 |||| 8 |||| 5,57 20 407
SemTab2022 HT-R2 Test 426 2 |||| 5 |||| 2,53 4 |||| 8 |||| 5,56 1 829
Train 9 917 2 |||| 4 |||| 2,51 3 |||| 11 |||| 5,65 64 542
SemTab2023 WikidataTables-R1 Test 500 2 |||| 4 |||| 2,46 3 |||| 11 |||| 6,95 4 247
SemTab2021 BioDiv-red Test† 11 1 |||| 26 |||| 17,45 26 |||| 100 |||| 58,90 1 232
TURL-120k Train 13 061 1 |||| 43 |||| 5,43 1 |||| 624 |||| 12,98 120 000
TURL-2k-red Test 1 295 1 |||| 14 |||| 1,03 6 |||| 257 |||| 32,95 1 801
4. Study Set-up
[…]
To maintain a good balance between speed, coverage, and diversity, we retrieve and inject at most $50$ candidates for every mention, eventually adding the correct one.
Our evaluation has the main objective of evaluating the selected approaches on datasets not considered in the original experiments, considering especially the evaluation of TURL and TableLlama on STI-derived datasets and of Alligator and Dagobah (with some limitations) on the datasets used to evaluate the first two approaches. Therefore, we first discuss the datasets used in our study (Section 4.1), and the evaluation protocol with its associated objectives (Section 4.1). Then, we provide details about ER (Section 4.3) and the training of the models (Section 4.4).
4.1. Datasets
The datasets considered in this work come from different sources and contain information about different domains. In particular, we have selected the following:
  • SemTab2021 - R3 (BioDiv) (abdelmageed2021biodivtab, ): the BioDiv dataset is a domain-specific benchmark comprising $50$ tables from biodiversity research extracted from original tabular data sources; annotations have been manually curated.
  • SemTab2022 - R2 (2T) (Cutrona2020, ): the dataset consists of a set of high-quality manually-curated tables with non-obviously linkable cells, i.e., where mentions are ambiguous names, typos, and misspelled entity names;/
  • SemTab2022 - R1 & R2 (HardTables (HT)) (abdelmageed2022results, ): datasets with tables generated using SPARQL queries (Jimenez2020, ). The datasets used from HardTables 2022 are round 1 (R1) and round 2 (R2);
  • SemTab2023 - R1 (WikidataTables) (hassanzadeh2023results, ): datasets with tables generated using SPARQL queries for creating realisticlooking tables. The dataset includes Test and Validation tables, yet we exclusively employ the Validation tables due to Gold Standard (GS) not being provided;
  • TURL (Deng2022, ): the authors of TURL developed the TURL dataset using the extensive WikiTable corpus, a rich compilation of tables from Wikipedia. This dataset includes table metadata such as the table name, caption, and column headers. This dataset has been sub-sampled by TableLlama’s authors to create a smaller version containing exactly 2K mentions, called TURL-2K, and used as the test set.
Table 3 reports the statistics of each dataset used for testing and fine-tuning the selected approaches (†indicates the datasets that have undergone a reduction). The SemTab datasets were already split in train and test except for BioDiv 2021, which has been only used during the testing phase along with the TURL-2K. To create a unified and common experimental setting, each dataset has been modified to contain the same set of mentions: datasets for TableLlama were first created by generating prompts for each mention while filtering out all prompts that exceeded TableLlama’s context length (which is equal to $8192$ tokens). For this reason, we have renamed BioDiv, 2T and TURL-2K into BioDiv-red, 2T-red and TURL-2K-red. To reduce the number of tokens generated for each prompt, we have taken some precautions: i) the description separator has been reduced from [DESCRIPTION] to [DESC], and ii) the row separator ([SEP]) has been deleted. Then, TURL datasets and the ground truths for Alligator and Dagobah were created using the same mentions as TableLlama.
Table 4. Test datasets denomination, as specified in Section 4.2. For every dataset, we report its source, the domain of the entities contained, and the characterization as In-Domain (IN), Out-Of-Domain (OOD), and Moderately-Out-Of-Domain (MOOD) for each of the pre-trained models (see Table 1 for the pre-training datasets). {Dataset}-red means that the specific dataset has been reduced as explained in Section 4.1.
Approaches
Dataset Source Entities Domain Alligator TURL TableLlama
SemTab2022 2T-red Graph Queries / Web Tables Cross IN MOOD MOOD
SemTab2022 HT-R1 Graph Queries Cross IN MOOD MOOD
SemTab2022 HT-R2 Graph Queries Cross IN MOOD MOOD
SemTab2023 WikidataTables-R1 Graph Queries Cross IN MOOD MOOD
SemTab2021 BioDiv-red Biodiversity Tables Specific OOD OOD OOD
TURL-120k Web Tables Cross MOOD IN IN
TURL-2K-red Web Tables Cross MOOD IN IN
4.2. Distribution-aware Experimental Objectives
The datasets used for training or evaluating ED are generated from different sources, contain tables of different sizes, and hold information about disparate domains. We assume each dataset is associated with a data distribution generating it (Jimenez2019, ; Jimenez2020, ; Cutrona2020, ; abdelmageed2021biodivtab, ; hassanzadeh2023results, ; Deng2022, ).
We introduce the concepts of “in-domain”, “out-of-domain” and “moderately-out-of-domain” settings related to the evaluation of a particular approach on a test set, specialising a distinction between “in-domain” and “out-of-domain” used in (Zhang2023, ). These denominations depend on the source the data has been generated from and the domain(s) it holds the information about. In particular, we define:
  • in-domain” (IN): a test set $Y$ for an approach $A$ is considered “in-domain” for $A$ if $A$ has been trained on a dataset $X$ generated from the same data source and covering similar domain(s) as $Y$;
  • out-of-domain” (OOD): a test set $Y$ for an approach $A$ is considered “out-of-domain” if $A$ has been trained on a set of data $X$ generated from a different data source and covering different domain(s) as $Y$;
  • moderately-out-of-domain” (MOOD): a test set $Y$ for an approach $A$ is considered “moderately-out-of-domain” for $A$ if $A$ has been trained on a set of data $X$ generated from the same data source but covering different domain(s) as $Y$ or vice versa, i.e., if $A$ has been trained on a set of data $X$ generated from a different source but covering similar domain(s) as $X$. This covers a setting such as the evaluation of TableLlama, pre-trained on the TURL dataset (Deng2022, ), on the STI-derived test set 2T: the two datasets contain different tables but cover cross-domain information linked to Wikidata in a quite similar way.
Given the definitions above, the evaluation procedure has been divided into two steps
  1. (1)
    We assess the performance of the considered approaches on the test data defined in Section 4.1 and detailed in Table 3, without further fine-tuning. The heterogeneity of the test data implies all the algorithms are tested against “in-domain”, “moderately-out-of-domain” and “out-of-domain” data.
  2. (2)
    We fine-tune TURL, TableLlama and Alligator on the train splits from their MOOD data, i.e., on SemTab2022 HT-R1, SemTab2022 HT-R2, SemTab2023 WikidataTables-R1 and SemTab2022 2T-red for TURL and TableLlama, and on TURL120k [^5] TURL-120k has been created starting from the original TURL dataset (Deng2022, ) by sub-sampling 13’061 tables containing 120’000 mentions ($\approx$ 125’000 is the number of mentions used to fine-tune TableLlama and TURL on their MOOD data combined). for Alligator; finally we test the fine-tuned models on our test data.
The main objectives of our analysis can be summarised as follows:
  • test the capability of pre-trained approaches on tables from different datasets, which we expect to be representative of different data distributions;
  • test the generalisation capability of both feature-based ML and LLM-based approaches after fine-tuning in MOOD settings;
  • compare the approaches in inference time and occupied memory to get insights about their applicability on application domains where processing of large tables may be required (e.g., enrichment of business data).
  • identify strengths and weaknesses of LLM-based approaches on the ED task with ablation studies.
4.3. Candidate entity retrieval with LamAPI
The candidates for the mentions contained in the TURL dataset (Deng2022, ), along with its sub-sampled version TURL-2K (Zhang2023, ), were retrieved through the Wikidata-Lookup-Service [^6] https://www.wikidata.org/w/api.php?action=wbsearchentities&search=Obama&language=en&limit=50&format=json, which retrieves at most $50$ candidates for the mention “Obama”, which is known to have a low coverage w.r.t. other ERs (Avogadro2022, ). For this reason, we researched to identify a state-of-the-art approach/tool specific to the ER. The final choice fell on LamAPI, an ER system developed to query and filter entities in a KG by applying complex string-matching algorithms. In LamAPI, an ElasticSearch777www.elastic.co index of the Wikidata dump (v. 01072024) has been constructed, leveraging an engine designed to search and analyse extensive data volumes in nearly real-time swiftly. The choice to use LamAPI as an ER system is based on its availability and performance compared to other available systems (e.g., Wikidata Lookup) (Avogadro2022, ). To further validate the choice of LamAPI we have computed the number of mentions with $K$ candidates for the TURL-2K-red dataset: the original TURL-2K dataset has almost $600$ mentions with $1$ candidate (the correct one, with $969$ mentions ($\approx 48\%$) with at most 5 candidates included the correct one). For all those mentions the Wikidata-Lookup service fails to retrieve something meaningful. On the contrary LamAPI retrieves for $1650$ mentions ($\approx 91\%$) in our sub-sampled dataset at least $45$ candidates. In particular, for the TURL-2K-red dataset, the coverage (i.e., how many times the correct candidate is retrieved by the ER system over the total number of mentions to cover) of LamAPI and Wikidata-Lookup is $88.17\%$ and $71.75\%$ respectively. For all these reasons we decided to replace the candidates extracted from LamAPI to build a new version of TURL-2K-red dataset which we called TURL-2K-red-LamAPI.
4.4. Training and implementation details
Pre-training and model usage. For TURL we first replicated the pre-training with the same hyperparameters as specified by the authors in (Deng2022, ) but in a distributed setting on 4 80GB-A100 GPUs, following the findings in (goyal2018accurate, ), then we fine-tuned it with the default hyper-parameters, matching the CEA results in the original paper. We use the open-source version of TableLlama made available on HuggingFace [^8] https://huggingface.co/osunlp/TableLlama.
Figure 2. Per-approach average accuracy, elapsed time (min), and occupied memory (GB). The time reported is the elapsed time to link a mention, without considering the ER time. The occupied memory is the occupied GPU memory for TURL and TableLlama, while is the “Virtual Memory Size” for Alligator and Dagobah. {Approach}-FT represents the respective approach fine-tuned in MOOD setting (Sec. 4.2). “O.O.M.” stands for Out-Of-Memory.
[…]
Dataset TURL TableLlama Alligator Dagobah GPT-4o GPT-4o-mini TURL-FT TableLlama-FT Alligator-FT
[…]
SemTab2021 BioDiv-red 0.8109 OOD 0.9513 OOD 0.5674 OOD N.A. OOD 0.9748 UNK 0.9497 UNK 0.6347 OOD 0.9610 OOD 0.8547 OOD
[…]
Alligator is pre-trained on different SemTab datasets before 2022 as in the original paper [^9] As reported in Table 1, the 2T dataset is one of the datasets used to pre-train Alligator. Even though we have 2T in our test data, those two datasets come from different rounds and years of the SemTab challenge.. We refer to Table 1 for details about the datasets used for pre-train. Dagobah is run with the default hyperparameters as specified in the corresponding GitHub repository [^10] https://github.com/Orange-OpenSource/Table-Annotation.
[…]
Evaluation and inference. Following our primary objective, all selected approaches were evaluated on the same test data (refer to Table 3). Both TableLlama and TURL were run with a batch size of 1, though this setting has distinct interpretations: for TURL, it means processing one table at a time, while for TableLlama, it means processing a single mention at a time. The GPT-4o and GPT-4o-mini models were also run with TableLlama-style prompts, meaning they process one mention at a time. In contrast, Alligator was run with its default configuration, using 8 parallel processes, each handling up to 50 rows per table. Lastly, Dagobah was executed with its default hyperparameters, as specified in its official GitHub repository.
Technical infrastructure. Both test and fine-tuning for TableLlama and TURL has run on the Azure NC24ads_a100_v4; Alligator on an Intel(R) Xeon(R) Gold 6238R CPU @ 2.20GHz with $64$ cores and $96$GB of RAM, while Dagobah on an Intel(R) Xeon(R) CPU E5-2650 0 @ 2.00GHz with $32$ cores and $96$GB of RAM.
5. Results and discussion
We first focus on the main results, then we discuss evidence from ablation studies.
5.1. Main results
Figure 2 reports the per-approach average accuracy, with the average computed over the test datasets (cfr. Table 3), while Table 5 reports the detailed performances achieved by the four different approaches on the test data. We observe that Alligator excels on HT-R1, HT-R2 and WikidataTablesR1 SemTab datasets, it performs poorly on both BioDiv-red and TURL-2K-red-LamAPI while performing discretely on 2T-red dataset w.r.t. TableLlama. The opposite is observed for TableLlama and TURL, with TableLlama achieving generally higher accuracy than TURL. This trend can be explained by the fact that TURL, TableLlama and Alligator encode different inductive biases: features computed by Alligator aim to capture signals coming from both the underline KG and the syntactical relatedness between a mention and its candidates; GTUM models like TURL and TableLlama on the other hand capture more semantic and high-level features, lacking the direct interaction with the KG. Moreover, the datasets they excel on are the ones considered in-domain (IN). Surprisingly, both TURL and TableLlama excel on BioDiv, even though it’s considered OOD for both approaches: we hypothesise that the enormous and general pre-training knowledge these models retain explains this excellence. The poor performance achieved by TURL on all the SemTab data, considered MOOD for it, can be explained by the fact that TURL is already fine-tuned on web tables, which are generally coming from a different data distribution than the ones of the SemTab challenge. Interestingly, we observed that both TableLlama and TURL underperforms themselves on the TURL-2K-LamAPI dataset: we argue that the drop in performances, especially for TURL, is due to the increased number of candidates we have retrieved with LamAPI (cfr. the ablation study in section 5.2). Regarding the GPT family, GPT-4o is, on average, the best-performing model, with GPT-4o-mini that is on par with the MOOD fine-tuned version of Alligator.
[…]
Figure 3. Per-approach accuracy with respect to data typology, as defined in Section 4.2. The average is computed over datasets sharing the same data typology.
MOOD fine-tuning. Table 5 reports also the performances of TURL, TableLlama and Alligator after MOOD fine-tuning. Thanks to the “continual finetuning” inspired by (ibrahim2024simple, ) TableLlama slightly increases its performances on (previously) MOOD and OOD (BioDiv-red) data, with no clue of suffering from catastrophic forgetting on its IN data (TURL-2K-red-LamAPI). TURL gains the most from the MOOD fine-tuning ($\approx$$+36\%$ on WikidataTables-R1) but suffers a severe drop in performance on both IN and OOD data, confirming the catastrophic forgetting suffered by models after fine-tuning (goodfellow2015catastrophicforgetting, ). Alligator achieves great performances on 2T-red, BioDiv-red and TURL-2K-red-LamAPI after MOOD fine-tuning, while suffering a contained drop in performance (apart from HT-R2, where it loses 10%), demonstrating a better generalization pattern than TURL. The average performance with respect to the data typology is depicted in Figure 3, which further demonstrates these findings.
Computational performance. Figure 2 also reports the average elapsed time (center) for the six tested approaches and the average occupied memory (right) by each approach (TURL and TableLlama’s occupied memory refers to the GPU memory, while the Alligator and Dagobah’s memory refers to the overall “Virtual Memory Size”). Note that time is on a logarithmic scale and is considered as the time employed by an approach to link a mention without considering the ER time [^11] The GPT-4o and GPT-4o-mini time is constrained by the limits imposed by OpenAI API rate limits. If TURL is the fastest and the lightest from both time and occupied memory perspectives and TableLlama is the slowest and heaviest, Alligator strikes as a good compromise w.r.t. to execution time and occupied memory. Dagobah, on the other hand, is completely out-of-bounds, especially if one is bounded by resource or budget constraints.
5.2. Ablation studies
[…]
Accuracies achieved by TableLlama and TURL w.r.t. the number of candidates. Refer to caption
Figure 4. Accuracies achieved by Alligator, GPT4o, TableLlama and TURL w.r.t. the number of candidates on the TURL-2K-red-LamAPI dataset.
To explain the drop in performances observed for TURL and TableLlama on the TURL-2K-red-LamAPI dataset, we have measured the accuracy w.r.t. the number of candidates per mention, with the intuition that the higher the number of candidates the lower the performances. Figure 4 reports the accuracy of TURL, TableLlama, GPT-4o and GPT-4o-mini given a different number of candidates per mention, also considering the correct one. Our intuitions are empirically confirmed by observing a drop in performance for both approaches, with a more severe one for TURL. This could happen because TURL aggregates the candidates of every mention in a table and passes them through a Transformer, increasing the context length to at most $O(NK)$, where $N$ is the number of mentions in a table and $K$ is the maximum number of candidates retrieved per mention. We ran an additional ablation study to measure the impact of the table’s metadata (e.g., the title of the Wikipedia page the table is found in, the section title, or the table caption to name a few) on the final accuracy achieved by TURL, TableLlama, GPT-4o, and GPT-4o-mini, testing on the TURL-2K-red-LamAPI dataset. We ran experiments with i) No-Meta, i.e., a setting where the page title, section title, and the table caption are removed; ii) No-Header, i.e., the table header is changed to $[\text{col}0,\text{col}1,...,\text{col}N]$ and iii) No-Meta-No-Header, i.e., the combination of both i) and ii). From Figure 5, we observe that TURL is heavily dependent on the table metadata, with a severe drop in performance when both metadata and header are removed. On the other hand, TableLlama and GPT-4o are almost unaffected by removing metadata from the prompt, indicating a higher generalisation capability and a greater focus on the context provided by the table itself rather than by the metadata.
5.3. Discussion
[…]
Accuracies achieved by TableLlama and TURL w.r.t. the presence of table’s metadata. Refer to caption
Figure 5. Accuracies achieved by GPT4o, TableLlama and TURL w.r.t. the presence of table’s metadata on the TURL-2K-red-LamAPI dataset.
Generative GTUM approaches with a high number of parameters have interesting properties in terms of accuracy, generalisation, and robustness to the number of candidates that are processed and the table metadata, especially on MOOD and OOD data, as found in Sections 5.1 and 5.2; however, specific STI approaches trained on in-domain data, and optionally fine-tuned on moderately-out-of-domain data, using a tiny fraction of these parameters can still be on par with generalistic approaches while being faster and occupying less memory. Generative approaches like GPT-4o and TableLlama are more promising than encoder-based ones in terms of performance, with a negligible risk of hallucinations; however, the comparison considered models of uncomparable size (7B vs 300M for TableLlama and TURL resp.), with TURL being the fastest approach among those that were tested. Regarding scalability, long context allows the encapsulation of large tables with a high number of candidates, but some tables cannot fit in the context, and when they do, the computation time and occupied memory increase proportionally to the table size: considering a small context table (few rows above and below the row containing the mention to be linked) to be fed to the model could be effective. On the budget side, training models of the size of TableLlama has still enormous costs (48 A100-80GB for 9 training days), so devising generative methods based on smaller LLMs, e.g., Phi (abdin2024phi3, ), could be an interesting research direction, although more sophisticated approaches may be needed to achieve the same level of generalisation and reliability. Moreover, Figure 6 shows the accuracy achieved by the tested approaches, both before and after MOOD fine-tuning, with respect to the average price [^12] For TableLlama, TURL and Alligator, prices are estimated using the Azure Pricing Calculator, selecting the NC24ads_a100_v4 for running GPU inference with TableLlama and TURL, and D48plds_v5 for running CPU inference and hosting Alligator. GPT-4o and GPT-4o-mini prices are estimated from the OpenAI prices page, considering a token as being equal to 4 characters. paid to host machines to run inference on all the test data. The figure shows how GPT-4o, while being the best-performing model on average, is also the most expensive one; GPT-4o-mini, TableLlama, TableLlama-FT, and Alligator-FT are comparable in terms of accuracy and price paid, although TableLlama costs is slightly higher. These results suggest that generative GTUM approaches are best suited when the tables to be annotated contain rich, highly semantic textual data. GPT-4o is the most accurate among these methods but also the most expensive. GPT-4o-mini offers a fair compromise between accuracy, cost, and computational demand. On the other hand, Alligator is more appropriate when signals from the knowledge graph (KG) have a significant impact on the final accuracy, especially when time and budget are limited. When fine-tuned, Alligator performs comparably to GPT-4o-mini and TableLlama in terms of accuracy with a reduced cost.
[…]
Figure 6. Per-approach accuracy over price, on average. The price is computed given the average time spent (in minutes) to run the evaluation over all the test data.
Finally, TableLlama and TURL are not NIL aware and TableLlama and GPT-4o, in particular, cannot return confidence scores associated with cell annotations, which may be useful when using these approaches to support revision: devising methods to estimate the uncertainty of labels computed by generative models may be an interesting research direction.
6. Conclusions and Future Works
In this study, we addressed the lack of comprehensive comparative analyses among different approaches to Semantic Table Interpretation (STI), focusing specifically on the Cell Entity Annotation (CEA) task. We selected four representative methods—including large generative Large Language Models (LLMs) and specialized STI models—and evaluated them under common conditions. Our evaluation spanned multiple dimensions: accuracy, generalisation, execution time, memory usage, and cost. We tested these methods across various scenarios, i.e., “in-domain”, “out-of-domain” and “moderately out-of-domain”. Our experiments suggest that approaches based on a large generative LLM excels in accuracy and generalisation, as demonstrated by the results on MOOD and OOD data (cfr. Table 5 and Figure 3), at the price of an increased execution and training time. TURL, an encoder-only model based on TinyBERT, is the most efficient, but lacks on generalisation capabilities. Moreover, our experiments suggest that specific STI models like Alligator, despite their limited number of parameters, can still be a viable option when resources and budget are limited, with performance improvements obtained after fine-tuning. Finally, GPT-4o, while being the most accurate, is also the most expensive, with GPT-4o-mini that represents a better trade-off with respect to accuracy and cost. Future works include the possibility to train a smaller and cheaper LLM-based table model, e.g., Phi (abdin2024phi3, ), while enabling also the handling of NIL entities and a score associated with cell annotations. Additionally, we aim to augment LLM-based approaches with advanced prompting techniques like In-Context Learning or Chain-of-Thought reasoning. This enhancement could leverage the inherent reasoning capabilities of LLMs and facilitate human-in-the-loop interactions to guide the model’s behavior. Another potential direction is to modify models like TURL and TableLlama to operate at the cell level rather than the table level. This change could reduce memory consumption and allow for the application of standard data augmentation techniques, potentially improving both performance and efficiency.
7. Acknowledgments
This work has been partially funded by the European innovation action enRichMyData (HE 101070284) and the Italian PRIN project Discount Quality for Responsible Data Science: Human-in- the-Loop for Quality Data (202248FWFS) funded by the European Community - Next Generation EU.
AI
Artificial Intelligence
BOW
Bag-Of-Words
BKG
Background Knowledge Graph
CTA
Column-Type Annotation
CEA
Cell-Entity Annotation
CPA
Columns-Property Annotation
CRF
Conditional Random Field
DL
Deep Learning
EL
Entity Linking
ED
Entity Disambiguation
ER
Entity Retrieval
GUI
Graphical User Interface
GDBT
Gradient Boosted Decision Tree classification model
GS
Gold Standard
GSs
Gold Standards
GTUM
Generalistic Table Understanding and Manipulation
KG
Knowledge Graph
KGs
Knowledge Graphs
KB
Knowledge Base
KBs
Knowledge Bases
LOD
Linked Open Data
CNEA
Cell-New Entity Annotation
NLP
Natural Language Processing
Regex
Regular Expressions
RDF
Resource Description Framework
RML
RDF Mapping Language
RDFS
Resource Description Framework Schema
SOTA
state-of-the-art
STI
Semantic Table Interpretation
SVM
Support Vector Machine
UI
User Interface
IR
Information Retrieval
ML
Machine Learning
NE
Named Entity
NER
Named Entity Recognition
NEL
Named Entity Linking
NIL
Not In Lexicon
NN
Neural Network
LIT
Literal
CRF
Conditional Random Field
CNN
Convolutional Neural Network
RNN
Recurrent Neural Network
HNN
Hybrid Neural Network
LLM
Large Language Model
LLMs
Large Language Models
LMs
Language Models
LM
Language Model
HDT
Hybrid Decision Tree
PFSMs
Probabilistic Finite-State Machines
PGM
Probabilistic Graphical Model
GFT
Google Fusion Tables
PoI
Points of Interest